# A Computational Tsirelson’s Theorem for the Value of Compiled XOR Games

### v1

@@ 1 Introduction
In a nonlocal game, a polynomially-bounded classical verifier interacts with multiple non-communicating provers. A set of predetermined rules governs whether the provers win or lose the game. The classical value of a nonlocal game is the highest winning probability that any classical (or unentangled) provers can achieve, whereas the quantum value is the highest winning probability that any quantum (potentially entangled) provers can achieve. Although the study of nonlocal games originated with Bell’s theorem [Bel64], they have become central in the study of quantum complexity theory, through their connection with multiprover interactive proofs [CHTW04]. This connection has led to many protocols for efficiently certifying classical and quantum computations in the two-prover setting [RUV13, CGJV20, Gri17, ${}^{+}$21].
It is interesting to ask whether results in the nonlocal game model can be translated to a model where a verifier interacts with a single prover. The single-prover setting is both theoretically appealing and practically motivated, since a non-communication assumption between provers can be challenging to enforce. Recently, Kalai et al. [KLVY23] proposed a generic procedure that compiles any $k$-player one-round nonlocal game into a single-prover protocol, assuming the existence of quantum homomorphic encryption [Mah20, Bra18]. In the same work, the authors showed that the compilation process preserves the classical value of the game. Shortly after, Natarajan and Zhang [NZ23] proved that the same compiler preserves the quantum value of the $2$-player CHSH game [CHSH69]. This was sufficient for them to achieve efficient delegation of BQP computations, but their work left open the question of whether the Kalai et al. compiler preserved the quantum value for a more general class of games.
A natural generalization of the CHSH game is the class of two-player XOR games [CHTW04], where the winning predicate depends on the XOR of the players’ single-bit answers. These games have a foundational place in the theory of nonlocal games thanks to a theorem of Tsirelson [Tsi80, Tsi87], which exactly characterizes their optimal quantum strategies in terms of a semidefinite program. This makes XOR games one of the few classes of nonlocal games where the optimal quantum strategies are well understood, and finding the quantum value is computationally tractable. As such, they are a convenient stepping stone towards our understanding of more general games.
@@ 1.1 Results
The negligible factors in Theorem 1.1 come from the fact that the compiled game assumes the existence of quantum homomorphic encryption, whose security is only computational. It is well-known that quantum homomorphic encryption can be constructed from the hardness of the standard learning with errors (LWE) problem [Mah18, Bra18], and thus our theorem holds up to the same assumption. A useful consequence of this result is that we can apply the extensive literature on XOR games to give new examples of well-behaved compiled games. For instance, we obtain the following new results as almost immediate corollaries of the proof of our main theorem.
  • Self-testing for all XOR games: Using a result from Slofstra [Slo11], we show that every compiled XOR game self-tests some properties of the provers’ measurement operators: in particular, for each XOR game, there is a set of relations between the Alice and Bob operators that are approximately satisfied for any near-optimal strategy for the compiled game (Theorem 3.6). As a special case, we recover the self-test for anticommutation in [NZ23].
  • Parallel repetition: Building on [CSUU08], we show that the quantum value of compiled parallel repeated XOR games satisfies perfect exponential decay in the number of parallel repetitions (Theorem 3.11). This is the first parallel repetition theorem for compiled games.
@@ 1.2 Technical Outline
We recall the basic setup of a nonlocal game. Before the interaction, Alice and Bob prepare a shared and possibly entangled state $|\psi\rangle$. Once this is done, they are separated and cannot communicate. The verifier samples a pair of challenges $(x,y)$ from a known distribution, and sends $x$ to Alice and $y$ to Bob. The players measure their respective quantum observables $A_{x}$ and $B_{y}$ on their respective shares of the state $|\psi\rangle$, and return outcomes $a,b$ to the verifier. The verifier decides whether the players win or lose by evaluating a deterministic predicate $V(x,y,a,b)\in\{0,1\}$.
The KLVY compiler converts any nonlocal game into an interaction between the verifier and a single computationally-bounded prover, the latter of which is asked to sequentially play the roles of Alice and then Bob. The setup of the compiled game is as follows:
@@
    The prover prepares an initial state $|\psi\rangle$. There is no assumption that this state lives in a bipartite Hilbert space. Then the verifier generates a key pair for a quantum fully homomorphic encryption scheme, and sends the public key to the prover.
@@
    The verifier samples a pair of challenges $(x,y)$ from a known distribution. It sends the encryption $c=\mathrm{Enc}(x)$ of the first challenge to the prover. The prover, first acting as Alice, measures the state and returns the outcome $\alpha$. In the honest case, this measurement is the homomorphic execution of the observable $A_{x}$.
@@
    The verifier sends $y$ to the prover in the clear. The prover, now acting as Bob, measures an observable $B_{y}$ on its residual state and returns the outcome $b$. There is no assumption that this observable commutes with the earlier measurement.
@@
It is easy to show that the compiled value is always at least the nonlocal value, as any nonlocal strategy can be implemented under the homomorphic encryption. It is far less obvious how to establish upper bounds on the compiled value. This is because all known general techniques for upper-bounding the value of a nonlocal game make essential use of the spatial separation between the two provers. Nevertheless, we are able to show that the two specific techniques used by [NZ23] to analyze the compiled CHSH game both generalize to all compiled XOR games. We believe the two approaches are complementary and improve our understanding of compiled nonlocal games from dual perspectives.
@@
Our first approach generalizes the “macroscopic locality” proof of soundness of [NZ23] for the CHSH game, which is based on arguments contained in [Roh14]. One can use the IND-CPA security of the homomorphic encryption to upper-bound the maximum possible probability with which any efficiently measurable Bob operator can guess the value of Alice’s question $x$. Intuitively, this allows us to establish that Bob’s strategy cannot depend, in any detectable way, on Alice’s question. Applying Jensen’s inequality to Bob’s operator, we can then bound, up to a negligible function, the bias of every quantum strategy for the compiled XOR game, by the bias of some vector strategy for the nonlocal XOR game. These vector strategies can be thought of as a vector solution to the semidefinite program which computes the bias of the XOR game. By Tsirelson’s theorem, every such vector strategy gives rise to a quantum strategy for the nonlocal XOR game, yielding the desired bound on the winning probability for the compiled game in terms of the optimal nonlocal value of the original XOR game. As a corollary, we obtain a one-sided rigidity or “self-testing” statement for any compiled XOR game: for any near-optimal strategy, it must hold that $A_{x}|\psi\rangle\approx\sum_{y}c_{xy}B_{y}A_{x}|\psi\rangle$ for explicit coefficients $c_{xy}$.
@@
Another (more abstract but much more general) approach to upper-bounding the quantum value of a nonlocal game is to associate with a game a formal polynomial $p$, whose variables are identified with the observables $A_{x}$ and $B_{y}$, such that $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ is precisely the winning probability of the strategy. (For example, in the case of CHSH, the polynomial is $p(A_{x},B_{y})=\frac{1}{2}+\frac{1}{8}(A_{0}B_{0}+A_{0}B_{1}+A_{1}B_{0}-A_{1}B_{1})$, and it can be easily checked that $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ captures the winning probability of the strategy ($(\{A_{x}\},\{B_{y}\},|\psi\rangle)$ in CHSH.) To prove that the (nonlocal) quantum value of a game is at most $\omega$, it then suffices to give a sum-of-squares (SOS) certificate that $\omega-p(A_{x},B_{y})$ is a positive operator whenever $p$ is evaluated on a setting of the variables that corresponds to valid observables. Such a certificate consists of a decomposition of $\omega-p(A_{x},B_{y})$ as a positive operator
@@
where the $q_{i}$’s and $g$ are polynomials, and $g$ is a polynomial that evaluates to zero whenever the variables are valid observables. In particular, $g$ lies in the ideal generated by a certain collection of polynomial constraints on $A_{x}$ and $B_{y}$, including the commutator constraint $[A_{x},B_{y}]=0$ for all $x,y$, which formally represents the spatial separation of the players.
The reason why such an SOS certificate suffices to upper bound the game value in the nonlocal case is that 1) taking the expectation $\langle\psi|\omega-p(A_{x},B_{y})|\psi\rangle$ maps positive operators to positive numbers, and 2) $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ is equal to the winning probability of strategy ($(\{A_{x},B_{y}\},|\psi\rangle)$. Property 1 gives that evaluating the expectation on $\omega-p(A_{x},B_{y})$ yields a positive number, and property 2 says that this positive number is equal to $\omega-p_{win}$, where $p_{win}$ is the winning probability of the strategy ($(\{A_{x},B_{y}\},|\psi\rangle)$. In the nonlocal case, therefore, taking the expectation plays the role of what is often known as a “pseudo-expectation” in the SOS literature.
In general, an SOS certificate of the aforementioned form does not automatically yield an upper bound on the compiled value. The reason is that there is no longer an obvious choice of pseudo-expectation operator which will satisfy the requisite properties. In particular, the game value can no longer be described by a sum of quantities that look like $\langle\psi|A_{x}B_{y}|\psi\rangle$, and also we no longer have the guarantee $[A_{x},B_{y}]=0$ which spatial separation afforded to us before, meaning that the $g$ polynomial may not evaluate to the zero operator on all “valid” strategies for the compiled game. In [NZ23], a choice of pseudo-expectation operator was carefully defined so that the pseudo-expectation of $g(A_{x},B_{y})$ was zero by fiat, and (weaker versions of) properties 1 and 2 could be argued by reduction to cryptographic security properties. However, the fact that the measurement operators used by Alice in the compiled game return encrypted outcomes was an obstacle to carrying out these cryptographic reductions because cryptographic security only holds with respect to efficient distinguishers. In [NZ23], this obstacle was sidestepped for the specific case of the CHSH game, because the SOS certificate for this game has a particularly simple dependence on the $A_{x}$ operators. Here we show, building on a work of Slofstra [Slo11], that any XOR game has a “nice” SOS certificate (in which each $q_{i}$ has degree $1$ and depends on at most one $A_{x}$ variable) and that SOS certificates of such a nice form yield bounds on the compiled value as well as on the nonlocal value.
@@ 1.3 Further directions
The most immediate question raised by our work is to determine how far our techniques can extend beyond XOR games: for instance, can we find “nice” SOS decompositions, of higher degree, for all games? Another more conceptual question is to determine the “correct” notion of rigidity in the compiled setting: unlike in the nonlocal case, our results are not symmetric between between Alice and Bob.
@@ 2.1 Cryptography
We adopt many definitions from [NZ23].
@@ 2.2 Non-Local Games
A quantum (tensor) strategy $S$ for a nonlocal game $\mathcal{G}=(I_{A},I_{B},O_{A},O_{B},\pi,V)$ is a tuple $S=(\mathcal{H}_{A},\mathcal{H}_{B},|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$, consisting of finite-dimensionals Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, a bipartite state $|\psi\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, PVMs $\{A_{xa}\}_{a\in O_{A}}$ acting on $\mathcal{H}_{A}$ for each $x\in I_{A}$ for Alice and PVMs $\{B_{yb}\}_{b\in O_{B}}$ acting on $\mathcal{H}_{B}$ for each $y\in I_{B}$ for Bob. Often we will drop the Hilbert spaces, and just write $S=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$.
@@
We can also define the commuting operator (also known as the quantum computing) value of a nonlocal game $\omega_{qc}^{*}(\mathcal{G})=\sup_{S}\omega_{qc}(\mathcal{S},\mathcal{G})$ to be the supremum over all commuting operator strategies $S$ compatible with $\mathcal{G}$. It is not hard to see that every quantum (tensor) strategy is a commuting operator strategy. The converse holds if we restrict ou commuting operator strategies to be finite-dimensional (i.e. $\mathcal{H}$ is finite-dimensional). However, there are examples of nonlocal games $\mathcal{G}$ for which $\omega^{*}_{qc}(\mathcal{G})>\omega^{*}_{q}(\mathcal{G})$, see for example [Slo16].
@@ 3.1 The quantum value of compiled XOR games
$\sqrt{\sum_{s,y\in I_{B}}G_{xs}G_{xy}V_{sy}}=\sum_{y\in I_{B}}G_{xy}\langle u_{x}|v_{y}\rangle,$ (3.1)
@@ 4 The SOS approach to bounding the compiled value
In this section, we provide an alternative proof of Theorem 3.1 using an SOS approach.

### v2

@@ 1 Introduction
In a nonlocal game, a polynomially-bounded classical verifier interacts with multiple non-communicating provers. A set of predetermined rules governs whether the provers win or lose the game. The classical value of a nonlocal game is the highest winning probability that any classical (or unentangled) provers can achieve, whereas the quantum value is the highest winning probability that any quantum (potentially entangled) provers can achieve. Although the study of nonlocal games originated with Bell’s theorem [Bel64], they have become central in the study of quantum complexity theory through their connection with multiprover interactive proofs [CHTW04]. This connection has led to many protocols for efficiently certifying classical and quantum computations in the two-prover setting [RUV13, CGJV20, Gri17, JNV+21].
A central question in the nonlocal game paradigm is whether the techniques and results can be translated to a model where a verifier interacts with a single prover. The single-prover setting is theoretically appealing and practically motivated since a non-communication assumption between provers can be challenging to enforce. Recently, Kalai et al. [KLVY23] proposed a generic procedure that compiles any $k$-player one-round nonlocal game into a multi-round single-prover protocol, using a quantum homomorphic encryption scheme [Mah20, Bra18]. In the same work, the authors showed that the compilation process preserves the classical value of the nonlocal game. Shortly after, Natarajan and Zhang [NZ23] proved that the same compiler preserves the quantum value of the $2$-player CHSH game [CHSH69]. This was sufficient to achieve efficient delegation of BQP computations, however, their work left open the question of whether the Kalai et al. compiler preserved the quantum value for a more general class of games.
A natural generalization of the CHSH game is the class of two-player XOR games [CHTW04], where the winning predicate depends on the XOR of the players’ single-bit answers. These games have a foundational place in the theory of nonlocal games thanks to an influential theorem of Tsirelson [Tsi80, Tsi87], which exactly characterizes their optimal quantum strategies in terms of a semidefinite program. This makes XOR games one of the few classes of nonlocal games where the optimal quantum strategies are well understood, and finding the quantum value is computationally tractable. As such, they are a convenient stepping stone towards understanding more general families of nonlocal games.
@@ 1.1 Results
The negligible factors in Theorem 1.1 come from the fact that the compiled game assumes the existence of quantum homomorphic encryption, whose security is computational and depends on chosen parameters. It is well-known that quantum homomorphic encryption can be constructed from the hardness of the standard learning with errors (LWE) problem [Mah18, Bra18], and thus our theorem holds up to the same assumption. A useful consequence of this result is that we can apply the extensive literature on XOR games to give new examples of well-behaved compiled games. For instance, we obtain the following new results as almost immediate corollaries of the proof of our main theorem:
  • Self-testing for all XOR games Using a result from Slofstra [Slo11], we show that every compiled XOR game self-tests some properties of the provers’ measurement operators: in particular, for each XOR game there is a set of relations between the Alice and Bob operators that are approximately satisfied for any near-optimal strategy for the compiled game (Theorem 3.6). As a special case, we recover the self-test for anticommutation in [NZ23].
  • Parallel repetition Building on [CSUU08], we show that the quantum value of compiled parallel repeated XOR games satisfies perfect exponential decay in the number of parallel repetitions (Theorem 3.11). This is the first parallel repetition theorem for compiled games.
@@ 1.2 Technical Outline
We recall the basic setup of a nonlocal game. Before the interaction, Alice and Bob prepare a shared and possibly entangled state $|\psi\rangle$. Once this is done, they are separated and cannot communicate. The verifier samples a pair of challenges (or questions) $(x,y)$ from a known distribution, and sends $x$ to Alice and $y$ to Bob. The players measure their respective quantum observables $A_{x}$ and $B_{y}$ on their respective shares of the state $|\psi\rangle$, and return outcomes $a,b$ to the verifier. The verifier decides whether the players win or lose by evaluating a deterministic predicate $V(x,y,a,b)\in\{0,1\}$.
The KLVY compiler converts any nonlocal game into an interaction between the verifier and a single computationally-bounded prover, the latter of which sequentially emulates the roles of Alice and Bob. The setup of the compiled game is as follows:
@@
    The prover prepares an initial state $|\psi\rangle$. There is no assumption that this state lives in a bipartite Hilbert space. Then the verifier generates a key pair for a quantum fully homomorphic encryption scheme and sends the public key to the prover.
@@
    The verifier samples a pair of challenges $(x,y)$ from a known distribution. It sends the encryption $c=\mathrm{Enc}(x)$ of the first challenge to the prover. The prover, acting first as Alice, measures the state and returns the outcome $\alpha$. In the honest case, this measurement is the homomorphic execution of measuring the observable $A_{x}$ with respect to $|\psi\rangle$.
@@
    The verifier sends $y$ to the prover in the clear (i.e. unencrypted). The prover, now acting as Bob, measures an observable $B_{y}$ on its residual state and returns the outcome $b$. There is no assumption that this observable commutes with the earlier “Alice” measurement.
@@
It is easy to show that the compiled value is always at least the nonlocal value, as any nonlocal strategy can be implemented using the homomorphic encryption scheme. It is far less obvious how to establish upper bounds on the compiled value. This is because all known general techniques for upper-bounding the value of a nonlocal game make essential use of the spatial separation between the two provers. Nevertheless, we are able to show that the two specific techniques used by [NZ23] to analyze the compiled CHSH game both generalize to all compiled XOR games. We believe the two approaches are complementary and improve our understanding of compiled nonlocal games from dual perspectives.
@@
Our first approach generalizes the “macroscopic locality” proof of soundness of [NZ23] for the CHSH game, which is based on arguments contained in [Roh14]. One can use the IND-CPA security of the homomorphic encryption to upper-bound the maximum possible probability with which any efficiently measurable Bob operator can guess the value of Alice’s question $x$. Intuitively, this allows us to establish that Bob’s strategy cannot depend, in any detectable way, on Alice’s question. Applying Jensen’s inequality to Bob’s operator, we can then bound, up to a negligible function, the bias of every quantum strategy for the compiled XOR game by the bias of some vector strategy for the nonlocal XOR game. These vector strategies can be thought of as a vector solution to the semidefinite program which computes the bias of the XOR game. By Tsirelson’s theorem, every such vector strategy gives rise to a quantum strategy for the nonlocal XOR game, yielding the desired bound on the winning probability for the compiled game in terms of the optimal nonlocal value of the original XOR game. As a corollary, we obtain a one-sided rigidity or “self-testing” statement for any compiled XOR game: For any near-optimal strategy, it must hold that $A_{x}|\psi\rangle\approx\sum_{y}c_{xy}B_{y}A_{x}|\psi\rangle$ for explicit coefficients $c_{xy}$.
@@
A more abstract approach to upper-bounding the quantum value of a nonlocal game is to associate with a game a formal polynomial $p$, whose variables are identified with the observables $A_{x}$ and $B_{y}$ so that $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ is precisely the winning probability of the strategy. (For example, in the case of CHSH, the polynomial is $p(A_{x},B_{y})=\frac{1}{2}+\frac{1}{8}(A_{0}B_{0}+A_{0}B_{1}+A_{1}B_{0}-A_{1}B_{1})$, and it can be easily verified that $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ captures the winning probability of the strategy ($(\{A_{x}\},\{B_{y}\},|\psi\rangle)$ for the CHSH game. Moreover, to prove that the quantum value of a game is at most $\omega$, it suffices to give a sum-of-squares (SOS) certificate demonstrating that $\omega-p(A_{x},B_{y})$ is a positive operator whenever $p$ is evaluated on variables that corresponds to valid observables. Such a certificate consists of a decomposition of $\omega-p(A_{x},B_{y})$ as a positive operator
@@
where the $q_{i}$’s and $g$ are polynomials, and $g$ is a polynomial that evaluates to zero whenever the variables are valid observables. In particular, $g$ lies in the ideal generated by a certain collection of polynomial constraints on the variables $A_{x}$ and $B_{y}$, including the commutator constraint $[A_{x},B_{y}]=0$ for all $x,y$, which formally represents the spatial separation of the players.
The reason why such an SOS certificate suffices to upper bound the game value in the nonlocal case is that (1) taking the expectation $\langle\psi|\omega-p(A_{x},B_{y})|\psi\rangle$ maps positive operators to positive numbers, and (2) $\langle\psi|p(A_{x},B_{y})|\psi\rangle$ is equal to the winning probability of strategy $S=(\{A_{x},B_{y}\},|\psi\rangle)$. Property (1) gives that evaluating the expectation on $\omega-p(A_{x},B_{y})$ yields a positive number, and property (2) says that this positive number is equal to $\omega-p_{win}$, where $p_{win}$ is the winning probability of the strategy $S=(\{A_{x},B_{y}\},|\psi\rangle)$. Therefore, in the nonlocal case, taking the expectation plays the role of what is often known as a “pseudo-expectation” in the SOS literature.
In general, an SOS certificate of the aforementioned form does not automatically yield an upper bound on the compiled value. The reason is that there is no longer an obvious choice of pseudo-expectation operator which will satisfy the requisite properties. In particular, the game value can no longer be described by a sum of quantities that look like $\langle\psi|A_{x}B_{y}|\psi\rangle$, as we no longer have the guarantee $[A_{x},B_{y}]=0$ afforded to us before by the spatial separation, meaning that the $g$ polynomial may not be zero when evaluated on the “valid” strategies for the compiled game. In [NZ23], a choice of pseudo-expectation operator was carefully defined so that the pseudo-expectation of $g(A_{x},B_{y})$ was zero by fiat, and (weaker versions of) properties (1) and (2) could be argued by reduction to cryptographic security properties. However, the fact that the measurement operators used by Alice in the compiled game return encrypted outcomes was an obstacle to carrying out these cryptographic reductions because the security only holds with respect to efficient distinguishers. In [NZ23], this obstacle was overcome in the CHSH game because the SOS certificate for this game has a particularly simple dependence on the $A_{x}$ observables. Building on a work of Slofstra [Slo11], we show that any XOR game has a “nice” SOS certificate (in which each $q_{i}$ has degree $1$ and depends on at most one $A_{x}$ variable) and that these “nice” SOS certificates yield bounds on the compiled value as well as the quantum value.
@@ 1.3 Further directions
The most immediate question raised by our work is to determine how far our techniques can extend beyond XOR games: for instance, can we find “nice” SOS decompositions, of higher degree, for all nonlocal games? Another more conceptual question is to determine the “correct” notion of rigidity in the compiled setting: unlike in the nonlocal case, our results are not symmetric between between Alice and Bob.
@@ 2.1 Cryptography
We adopt several definitions from [NZ23].
@@ 2.2 Non-Local Games
A quantum (tensor) strategy $S$ for a nonlocal game $\mathcal{G}=(I_{A},I_{B},O_{A},O_{B},\pi,V)$ is a tuple $S=(\mathcal{H}_{A},\mathcal{H}_{B},|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$, consisting of finite dimensionals Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, a bipartite state $|\psi\rangle\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, PVMs $\{A_{xa}\}_{a\in O_{A}}$ acting on $\mathcal{H}_{A}$ for each $x\in I_{A}$ for Alice and PVMs $\{B_{yb}\}_{b\in O_{B}}$ acting on $\mathcal{H}_{B}$ for each $y\in I_{B}$ for Bob. Often we will drop the Hilbert spaces, and just write $S=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$.
@@
We can also define the commuting operator (also known as the quantum commuting) value of a nonlocal game $\omega_{qc}^{*}(\mathcal{G})=\sup_{S}\omega_{qc}(\mathcal{S},\mathcal{G})$ to be the supremum over all commuting operator strategies $S$ compatible with $\mathcal{G}$. It is not hard to see that every quantum (tensor) strategy is a commuting operator strategy. The converse holds if we restrict our commuting operator strategies to be finite dimensional (i.e. $\mathcal{H}$ is finite dimensional). However, there are examples of nonlocal games $\mathcal{G}$ for which there is a perfect (wins with probability $1$) commuting operators strategy but no perfect tensor-product strategy, see for example [Slo16].
@@ 3.1 The quantum value of compiled XOR games
$\sqrt{\sum_{s,y\in I_{B}}G_{xs}G_{xy}V_{sy}}=\sum_{y\in I_{B}}G_{xy}\langle u_{x}|v_{y}\rangle.$ (3.1)
@@ 4 The SOS approach to bounding the compiled value
In this section, we provide an alternative proof of Theorem 3.1 using the SOS approach.